CH/DOC/93/03

Run Plans

First Run at the ANU

Author: B.R. Fulton

Date: 3 March 1993

Framework for Running at the ANU

There follows a first attempt at producing a framework for planning the transfer of our programme to ANU and the first period of running there. I have discussed it with a few people, but not had a chance yet to get everyones comments. Some parts may have to change in the light of availability etc. At this stage I have not been too specific about dates, but have tried to outline what activities need to be carried out in particular periods, and who is responsible for them. Once we have agreed this pattern, the next step will be to fix specific dates.

A few general points first. The pattern comprises four main periods,

1. dismantling and packaging the equipment at Daresbury,

2. unpacking and setting up the equipment at ANU,

3. learning to operate the accelerator and performing a few short tests of the setup with beam, and

4. the running preiod.

The personnel involved in each period will vary, some people will make their contribution in the early stages, others during the running preiod. All these are important contributions, and everyone involved will be considered part of the experiments even if they do not happen to be there for the actual beamtime.

During the visits it is hoped that we will all operate on the same renumeration rules. These will be that every member gets the same per-diem rate for food, phonecalls etc., while housing costs and car rentals (if needed) are paid centrally. Since we are all in things together we would not wish any different tratment. This will involve some cordination between the groups and the SERC office funding students trips.

Accommodation will not necessarily be the same for everyone. Since we will have people out from July until October (see timetable below) it is possible we may be able to obtain a house or apartment for “long stay” personnel. For others who are out for shorter periods over runs different accommodation will be appropriate (University guest rooms or hotels).

We hope to have a couple of SUN’s out at ANU while we are ther, so students in particular should be able to continue analysis, and of course email contact with the UK will be available.

The period of the main runs has been slipped back to October (maybe end September) to allow all students to attend the Summer School. The exact dates will probably not be fixed for a month or so. It is proposed that we take the same approach for the running period as is our practice on the NSF — that is to say the prime responsibility for getting a run setup and run is the responsibility of the relevant group, although they will of course call on others for help with the shift running. Furthermore, it is felt that having everyone out at th same time would not make sense, with some 15 people around it would be too many for everyone to contribute effectively. For this reason it is suggested the running be split into two sections with a 4–5 day gap between them. The order of running can be decided later, but the division suggested below is based mostly on my understanding of the availability of staff members.

Once we get to the next stage of planning it will be possible to fix dates for individuals contributions to the setup phases and for groups contributions to the running phase. Once this is done it is assumed individuals are committed to the periods agreed (setup phases) and that groups are responsible for having their members available for the running phase. Of course individuals may wish to take the opportunity for a vacation while they are there, and this can be arranged privately outside the times when committed to the work.

Timetable for the ANU

Phase 0

JSL and BRF will go out to ANU for a week at the end of March. During this visit we hope to cover details of the work needed at that end for the installation of the new mounts and feedthroughs in the chamber, the cabling to the control room and the siting of th elctronics and data acquisition. We will also try and settle spcific dates for the main running periods and for the test runs. We will also try to arrange some form of house or apartment rental for the people involved in phases 2 and 3. Before we go we will endeavor to talk with the relevant people involved in all the work going on to find out particular points we should follow up. We will be back from this trip in time to report to the collaboration meeting on 8 April, and in time for any necessary actions to be taken before equipmwnt is packaged up for shipment.

Phase 1

This will occur during the first 2–3 weeks of April at the NSF. It will involve dismantling all our electronics and data acquisition setup, labling and itemising all the units and preparing them for packaging by the shipping company. In addition, the main cable runs will be packaged up along with other items of hardware which we need to send (mounts, gas system, connecting cables, connectors manuals etc.).

This will be a big job and everyone is asked to be available to be called upon for help when needed. Doug has taken the inititive in co-ordinating contact with the shipping company and hopefully can take this responsibility. I will ask Steve and Roger to take responsibilty for itemising and marking all the other stuff, and perhaps Dick might be able to help them with the electronics/DA components. Vic and Bob will have to be involved with the computing and gas system side as well.

Doug has explored shipping and it looks like some 10 weeks will be needed between handing over the stuff and having it available in ANU. This means the stuff has to be ready at DL by 23 April at the very latest. That gives us about three weeks after the Glasgow conference! Our present run finishes 12 March, and we should be able to make a start then. However, we need to be careful in case there is a chance for another run at short notice, and also we have to leave the EM4 setup until after a possible Edinburgh run (later in March).

Phase 2

This will start at ANU in early July and will involve un-packing the equipment, getting the cables installed, making the changes to the chamber and setting up the electronics and data acquisition. The length of time this will take is hard to guess, and also depends on how many people we have there. My guess is that the main bulk of the work should be completed in about a month. However, there are likely to be some residual things left over to complete. These can be carried out by those staying on afterwards for phase 3 (see below)

The suggested personel for this period are BRF, GT, SJAH, AM and ML. The aim would be to have three people there at all times, with more at times of overlap. Each person would spend about three weeks (with the exception of SJAH who stays on for the next phase).

Hopefully we will be able to rent a house or apartment which should accommodate everyone during this preiod, and also those arriving later for phase 2.

BRF, GT and SJAH are agreed to making this period. AM and ML are happy as well, although funding has still to be arranged with SERC!

Phase 3

During August and September the aim is to have about three “long-stay” people out. Their job will be to complete any outstanding work on the setup. More importantly, they will spend as much time as they can sitting in on experiments so that they get to know the setup there, and most importantly learn how to operate the machine. These lucky people then become our resident experts who will have to show everyone else the ropes during the period of the main runs. Hence they will be staying through phase 4 as well. Sometime in September this group will have a couple of short (2–3 shift) runs on the machine. The main aim of these is to shake out any problems in the system, but it may also be an oportunity to test possible mods to the hybrids designed to improve efficiency (RAC will be looking into a few ideas he has over the spring and summer months) and to test a version of the new readout system (RH is hoping to have this available in time).

Personel for this period are suggested as SJAH, RPW and CDJ (can we put our trust in this team!). SJAH will already have been there for a month, and so will probably have a break during this period. The others will arrive during August sometime.

One thing I have not yet included in the plans, and which may be relevant to this period (and possbily the preceeding one) is the expertise of RH, VP and RAC. It may be that we will need their skills at periods during the setup phase or during the test runs.

Phase 4

The exact dates of the running period are not set, but it will probably start late September or early October. The specific experiments are not yet clear either, since in many cases this will be guided by the outcome of analysis currently in progress. At present I have anticipated experiments from each of the groups (with possibly two from “Birmingham” if Martin Freer arrives early). As I metioned earlier, it would be absurd to have some 13–14 people trying to contribute at the same time, so the suggestion is we split the running into two parts. Lets assume 2–3 runs of 4 days each, a break of 4–5 days and then another couple of runs of 4–5 days each (this is only a guide at present). One of the main influences on the order or runs is the availability of staff members, which in turn depends on the dates of different terms. My suggestion is that we plan to run the “Birmingham” and “Surrey” experiments in the first phase and the “York” and “Oxford” experiments in the second phase. Each group would be responsible for the planning and execution of “their” experiment, but would also have the members of the other group there to help. In addition, of course, there will be the three long-stay “experts(!)”. Under this senario each group would plan to arrive about a week before their run started to get aclimatised and “trained” by the “experts”. There is likely to be a period of a week when we would have a large number of people out (at the end of the first set of runs), but I don’t see how we can avoid this. It will make it very difficult in the control room, but hopefully the new group can be learning the ropes separately from the group finishing the experiment. I hope that the lab can cope with such a large number of visitors at the same time — it is imposing a lot on them.

As I noted at the start, this is a first go at the plans and there may be some things overlooked, or some things which won’t work. If anyone has any thoughts along these lines I’d appreciate comment. Please, however, don’t try to be too detailed at this stage (for example suggesting specific dates you can be there). At this stage it is important to get the overall framework sorted out. Once we know this is sound I will get in touch again and we can start to be specific. It probably doesn’t make sense to do this until after we get back from the trip to ANU at the end of this month, as many things could emerge from that which influence things. At the moment the only firm committment I think we need from people is to agree the responsibilities on them for phase 0 (the packaging), and for everyone to be available to help with the physical work during the first few weeks of April.

Brian